# Mixed-state geometric phases of coherent and squeezed spin states

### v1

@@ V Implications
Since the CSS and SSS have been realized in condensed matter or atomic, molecular, and optical systems [53, 54, 70, 71, 72], experimental generation and measurement of the mixed-state geometric phases may become feasible in the future. However, imposing the parallel-transport condition for the IGP or Uhlmann phase requires precise controls of the systems and the ancilla representing the environment, which can be challenging in laboratories. Nevertheless, the non-adiabatic IGP relaxes the constraint and becomes a prominent candidates for demonstrating interesting geometric effects of CSS and SSS at finite temperatures. We caution that the non-adiabatic IGP is buried in the total phase, and a systematic removal of the non-geometric contribution according to Eq. (19) is required to obtain the genuine geometric contribution.
On the other hand, the rapid development of quantum computers opens possibilities of simulating complex quantum systems. By using an ancilla qubit acting as the environment and coupling it to a system qubit, Ref. [73] shows that the Uhlmann phase of a spin-$1/2$ system can be simulated on a quantum computer by using two engineered Hamiltonians to respectively evolve the system and ancilla simultaneously. Since a spin-$j$ object may be decomposed as a collection of $2j$ spin-$1/2$ objects, one may use $2j$ qubits as a system of a spin-$j$ system coupled to an ancilla system of $2j$ qubits to realize the CSS or SSS coupled to the environment. It then follows a series of technical challenges to construct suitable engineered Hamiltonians for the system and ancilla in order to generate the corresponding IGP or Uhlmann phase when the corresponding parallel-transport condition is imposed during the evolution. After the targeted geometric phase is generated on a quantum computer, one may follow Ref. [73] to couple additional measurement qubits to the composite system and perform state tomography to extract the phase accumulated during the constrained evolution.
@@ VI Conclusion
We have derived the Uhlmann phase and the IGP along with its non-adiabatic generalization for the CSS and SSS. Selected examples are fully solved to demonstrate interesting properties of those mixed-state geometric phases. The Uhlmann phase is well-defined for both CSS and SSS due to its rigorous mathematical structure, but there may or many not be topological phase transitions manifested by discrete jumps of the Uhlmann phase as temperature varies. In contrast, the parallel-transport condition for the IGP does not have a general solution for the SSS although the IGP of the CSS exhibits discrete jumps. The non-adiabatic IGP forgoes the constraint and exhibits discrete jumps in the geometric part of the accumulated phase for the CSS. However, no finite-temperature jump is found in the non-adiabatic IGP for the $j=1$ SSS. Our formalisms provide the foundation for future theoretical and experimental investigations of mixed-state topological properties of quantum spin systems.
@@ ACKNOWLEDGEMENTS
H.G. was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904) and the National Natural Science Foundation of China (Grant No. 12074064). X.Y.H. was supported by the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2023ZB611). C.C.C. was supported by the National Science Foundation under Grant No. PHY-2310656.
@@ Appendix A Correspondence between Uhlmann phase and Berry phase of CSS as $T\rightarrow 0$
where $J_{z}|j,-j\rangle=-j|j,-j\rangle$ has been applied. This result precisely reduces to the Berry phase of the ground state $|j,-j\rangle$ [56]. More specifically, when $j=\frac{3}{2}$ and $\theta=\frac{\pi}{2}$, $|\zeta|=1$ and thus $\theta_{U}=\mathrm{i}\frac{3}{4}{\oint}_{\gamma}(-2\mathrm{i}\tan^{2}\frac{\theta}{2})\mathrm{d}\phi=3\pi\equiv\pi\mod 2\pi$, which confirms the result shown in Fig. 1 as $T\rightarrow 0$.

### v2

@@ V Implications
Since the CSS and SSS have been realized in condensed matter or atomic, molecular, and optical systems Estève et al. (2008); Gross et al. (2010); Huang et al. (2021); Byrnes (2024); Hosten et al. (2016), experimental generation and measurement of the mixed-state geometric phases may become feasible in the future. However, imposing the parallel-transport condition for the IGP or Uhlmann phase requires precise controls of the systems and the ancilla representing the environment, which can be challenging in laboratories. Nevertheless, the non-adiabatic IGP relaxes the constraint and becomes a prominent candidates for demonstrating interesting geometric effects of CSS and SSS at finite temperatures. We caution that the non-adiabatic IGP is buried in the total phase, and a systematic removal of the non-geometric contribution according to Eq. (22) is required to obtain the genuine geometric contribution.
On the other hand, the rapid development of quantum computers opens possibilities of simulating complex quantum systems. By using an ancilla qubit acting as the environment and coupling it to a system qubit, Ref. Viyuela et al. (2018) shows that the Uhlmann phase of a spin-$1/2$ system can be simulated on a quantum computer by using two engineered Hamiltonians to respectively evolve the system and ancilla simultaneously. Since a spin-$j$ object may be decomposed as a collection of $2j$ spin-$1/2$ objects, one may use $2j$ qubits as a system of a spin-$j$ system coupled to an ancilla system of $2j$ qubits to realize the CSS or SSS coupled to the environment. It then follows a series of technical challenges to construct suitable engineered Hamiltonians for the system and ancilla in order to generate the corresponding IGP or Uhlmann phase when the corresponding parallel-transport condition is imposed during the evolution. After the targeted geometric phase is generated on a quantum computer, one may follow Ref. Viyuela et al. (2018) to couple additional measurement qubits to the composite system and perform state tomography to extract the phase accumulated during the constrained evolution.
@@ VI Conclusion
We have derived the Uhlmann phase and the IGP along with its non-adiabatic generalization for the CSS and SSS. Selected examples are fully solved to demonstrate interesting properties of those mixed-state geometric phases. The Uhlmann phase is well-defined for both CSS and SSS due to its rigorous mathematical structure, but there may or many not be topological phase transitions manifested by discrete jumps of the Uhlmann phase as temperature varies. In contrast, the parallel-transport condition for the IGP only renders a trivial solution for the one-axis SSS although the IGP of the CSS exhibits discrete jumps. The non-adiabatic IGP forgoes the constraint and exhibits discrete jumps in the geometric part of the accumulated phase for the CSS. However, no finite-temperature jump is found in the non-adiabatic IGP for both one-axis and two-axis squeezing with $j=1$. Our formalisms provide the foundation for future theoretical and experimental investigations of mixed-state topological properties of quantum spin systems.
@@ ACKNOWLEDGEMENTS
H.G. was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904) and the National Natural Science Foundation of China (Grant No. 12074064). X.Y.H. was supported by the National Natural Science Foundation of China (Grant No. 12405008) and the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2023ZB611). C.C.C. was supported by the National Science Foundation under Grant No. PHY-2310656.
@@ Appendix A Correspondence between Uhlmann phase and Berry phase of CSS as $T\rightarrow 0$
where $J_{z}|j,-j\rangle=-j|j,-j\rangle$ has been applied. This result precisely reduces to the Berry phase of the ground state $|j,-j\rangle$ Chaturvedi et al. (1987). More specifically, when $j=\frac{3}{2}$ and $\theta=\frac{\pi}{2}$, $|\zeta|=1$ and thus $\theta_{U}=\mathrm{i}\frac{3}{4}{\oint}_{\gamma}(-2\mathrm{i}\tan^{2}\frac{\theta}{2})\mathrm{d}\phi=3\pi\equiv\pi\mod 2\pi$, which confirms the result shown in Fig. 1 as $T\rightarrow 0$.